Solve for C: 2cy - y = 2c - 1
Answer by nerdybill(7384)

2cy - y = 2c - 1
2cy = 2c - 1 + y
2cy - 2c = - 1 + y
c(2y - 2) = y-1
c = (y-1)/(2y - 2)